Beta of a portfolio. the beta of four stocks---- G, H ,I ,J---- are 0.41, 0.74, 1.14, and 1.59, respectively. what is the beta of a portfolio with the following weights in each asset?

                    weight G. weight H.   weight I.   weight J
portfolio 1.      25%        25%            25%           25%
portfolio 2.      30%        40%            20%           20%
portfolio 3.     10%         20%            40%           30%

what is the beta of all portfolios?
Answer #1

Beta of portfolio 1=25%*0.41+25%*0.74+25%*1.14+25%*1.59=0.97

Beta of portfolio 2=30%*0.41+40%*0.74+20%*1.14+20%*1.59=0.965

Beta of portfolio 3=10%*0.41+20%*0.74+40%*1.14+30%*1.59=1.122